2y - 1/6y =

This is from Introduction to College mATHEMATICS BY gARY d jONES AND wADI e mAHFOUD

Answer by checkley77(12844)

2y-(1/6)y=2Y-Y/6=(6*2Y-Y)/6=(12Y-Y)/6=11Y/6 ANSWER.
OR:
2Y-1/(6Y)=(2Y*6Y-1)/6Y=(12Y^2-1)/6 ANSWER.
OR:
(2Y-1)/6Y ANSWER.